Three liquids are at temperatures of 7 ◦C,
20◦C, and 34◦C, respectively. Equal masses of
the first two liquids are mixed, and the equilibrium temperature is 11◦C. Equal masses of
the second and third are then mixed, and the
equilibrium temperature is 22.6◦C.
Find the equilibrium temperature when
equal masses of the first and third are mixed.
Answer in units of ◦C.
1
Expert's answer
2020-04-06T09:20:35-0400

"m" - mass of each liquids (all masses are equal )

"C_1" - specific heat of the first third 

"C_2" - specific heat of the second  liquid

"C_3" - specific heat of the third liquid

Temperature of liquids: "T_1=7 \u25e6C,T_2=20\u25e6C, T_3=34\u25e6C"

Temperature of 1+2 liquids mix: "T_{12}=11^oC"

Temperature of 2+3 liquids mix: "T_{23}=22.6^oC"

Temperature of 1+3 liquids mix: "T_{13} - ?"


When the first two are mixed:

"m C₁ (T₁ − T_{12}) + m C₂ (T₂ − T_{12}) = 0 \\
C₁ (7− 11) + C₂ (20 − 11) = 0\\
4C_1=9C_2\\
C_1=2.25C_2"


When the second and therd are mixed:

"m C_2 (T_2 − T_{23}) + m C_3 (T_3 − T_{23}) = 0 \\
C_2 (20−22.6) + C₂ (34 −22.6) = 0\\
2.6C_2=11.4C_3\\
C_2=4.38C_3"


When the first and therd are mixed:

"m C_1 (T_1 − T_{13}) + m C_3 (T_3 − T_{13}) = 0 \\
C_1 (7−T_{13}) + C_3 (34 −T_{13}) = 0\\
C_1=2.25C_2=2.25(4.38C_3)=9.86C_3\\
9.86C_3 (7−T_{13})=-C_3(34 −T_{13})\\
9.86 (7−T_{13})=-(34 −T_{13})\\
T_{13}=9.5^oC"


"T_{13}=9.5^oC"